Find the magnitude of the charge if it has a Voltage of 13 V with a distance of 25 mm.


1
Expert's answer
2022-03-14T13:07:22-0400

"V=\frac{kq}{r}"

"q=\frac{Vr}{k}"

"q=\frac{13\times25\times10^{-3}}{9\times10^9}=3.61\times10^{-11}C"